OpenStudy (anonymous):

Math help please! A firecracker shoots up from a hill 140 feet high, with an initial speed of 100 feet per second. Using the formula H(t) = -16t2 + vt + s, approximately how long will it take the firecracker to hit the ground? (4 points) Five seconds Seven seconds Nine seconds 11 seconds

OpenStudy (anonymous):

@gina123 @gina123

OpenStudy (anonymous):

@Michele_Laino @ganeshie8

OpenStudy (anonymous):

i think the answer is 7 seconds

OpenStudy (anonymous):

Okay so here you have your equation H(t) = -16t2 + vt + s You are given the initial height = 140 = s And the initial speed = v = 100 ft/s 16 t^2 + 100 t + 140 = H(t) Since you want to find the time it took to hit the ground, you are looking for the x intercept. Set the equation equal to 0 then use the quadratic formula. 16 t^2 + 100 t + 140 = 0

OpenStudy (anonymous):

ok i understand that but im unsure of how to solve it

OpenStudy (anonymous):

Do you know the quadratic formula?

OpenStudy (anonymous):

yes, so i would just use that?!!!

OpenStudy (anonymous):

Yes. DO you see how the formula resembles a quadratic. The only difference is that it is a t and not an x

OpenStudy (anonymous):

You should get 6.53. That rounds to 7.

OpenStudy (anonymous):

-16 t^2 + 100 t + 140 = 0 I forgot the negative.

OpenStudy (anonymous):

ok yes i got that answer thanks so much!!!!!!! ;)

OpenStudy (anonymous):

No problem!!

OpenStudy (anonymous):

do you mind helping with one more?

OpenStudy (anonymous):

Darlene kicks a soccer ball off the ground and in the air, with an initial velocity of 34 feet per second. Using the formula H(t) = -16t2 + vt + s, what is the maximum height the soccer ball reaches? (4 points) 17.7 feet 18.1 feet 19.3 feet 20.2 feet Im getting 17.7 feet?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

height + velocity + gravity https://www.youtube.com/watch?v=K7b-da1o3BY

OpenStudy (anonymous):

Well what do you know? What is your initial velocity and initial height?

OpenStudy (anonymous):

the velocity is 34 feet and i believe the unital height is 0 feet

OpenStudy (anonymous):

assuming since the problem doesnt say

OpenStudy (anonymous):

for the height

OpenStudy (anonymous):

Yup. So plug them into the equation for v and s

OpenStudy (anonymous):

H(t) = -16t2 + 34t + 0

OpenStudy (anonymous):

yes so the answer would be 18.1 feet ??

OpenStudy (anonymous):

Yeah!

OpenStudy (anonymous):

Thanks so much for walking me through it!!!!!!!

OpenStudy (anonymous):

You're welcome (:

OpenStudy (anonymous):

Medal?